OpenStudy (anonymous):

Find f(x) and g(x) so the function can be expressed as y = f(g(x)) y = Seven divided by x squared. + 10

OpenStudy (anonymous):

@dan815

OpenStudy (zzr0ck3r):

\(f(g(x)) = \dfrac{7}{x^2+10}\)?

OpenStudy (zzr0ck3r):

@ivorypolnett

OpenStudy (zzr0ck3r):

or \(f(g(x)) = \dfrac{7}{x^2}+10\)

OpenStudy (anonymous):

the second one

OpenStudy (zzr0ck3r):

right I guess that was the point of your period. ok I will give you a hint for one way to do this \(f(g(x)) = f(\color{blue}{\frac{7}{x^2}})=\color{blue}{\frac{7}{x^2}}+10\)

OpenStudy (zzr0ck3r):

what do you think the function \(f\) needs to be for this to be a true statement?

OpenStudy (zzr0ck3r):

In other words, what function takes something and adds \(10\) to it?

OpenStudy (anonymous):

I'm honestly not sure.. im sorry

OpenStudy (anonymous):

@ivorypolnett I think a part of the question is missing. f(x) should be given before hand

OpenStudy (anonymous):

the problem doesn't give me f(x), im sorry @Harindu

OpenStudy (anonymous):

or even g(x)? then there are lots of answers :)